# Krijgslaan, 281-S9, 9000 Gent, Belgium

### v1

@@ 4.4.2 The oscillatory wormhole E/M gauge field background
The “Centaur” geometries of Anninos and Hofman Anninos:2017hhn , can in a sense be thought of as a finely tuned case of the wineglass wormholes, when the running scalar reaches a local minimum/maximum of the scalar potential with positive value and stops completely there. On the one hand, it seems that it is always possible to achieve this, for example if the potential $V(\phi)$ becomes exactly flat after a certain value of $\phi$. Unfortunately the presence of the magnetic flux/radiation, which is very important for supporting the throat at its minimum value $a_{\rm min}(\tau_{\rm min})$, prohibits also the smooth closure of the geometry at other Euclidean times.
@@ 5.1 Dirichlet for the scalar field/Dirichlet for the gauge field
Thus, we expect $\Delta\mathcal{S}_{D-C}\underset{H_{\rm UV}\to\infty}{\sim}S_{E,\rm rad}^{D}>0$ for large $H_{\rm UV}$, i.e., the simple wormhole being more dominant in the gravitational path integral. The complete numerical results are shown fig. 8  [^23] In the right panel, we also show the results imposing Neumann BCs for the gauge field. The results are qualitatively similar.. In a certain regime there exist two wormholes: one is always subdominant relative to the disconnected geometry and ceases to exist when reaching the critical value $(\kappa|\tilde{V}_{\rm AdS}|)^{1/2}H_{\rm UV}=2\sqrt{2}$. This is the bound found earlier for the coexistence of two simple wormholes, see fig. 2. Keeping the dominant saddle between these two wormholes, we observe an exchange of dominance between the disconnected geometry and the wormhole [^24] This was also observed in two dimensional string theory models that exhibit BH/wormhole transitions (two disconnected cigars vs. a simple Euclidean wormhole) Betzios:2023jco . revealing a first order phase transition similar to the Hawking-Page transition at $(\kappa|\tilde{V}_{\rm AdS}|)^{1/2}H_{\rm UV}\simeq 2.708$ Hawking:1982dh .
@@ 5.2 Neumann for the scalar field/Dirichlet for the gauge field
The second line corresponds precisely to the counterterm and the boundary term. The counterterm is computed in Appendix A.1, and the boundary term vanishes on-shell. This is because, when $c_{\rm AdS}\neq 2$, the mode $\tilde{\phi}_{+}$ (proportional to the source using the Neumann boundary condition) is zero. Note that, in contrast to the preceding section, $c_{\rm AdS}$ is now related to a VEV and thus free to vary, resulting in a four-parameter family which we parametrize in terms of $H_{\rm UV},\,\tilde{V}_{\rm AdS},\,A_{\rm dS},\,A_{\rm AdS}$. After sending the cut-offs to infinity, the difference between the on-shell actions remains finite
@@ 6 Discussion
    Semiclassical saddle competition in the gravitational path integral and Cosmological ramifications: By explicitly evaluating the on-shell action differences, we identified the parameter regions in which different saddles dominate the gravitational path integral. As we explain in section 2, the connected two boundary saddles can be used to estimate ratios of probabilities for different cosmological outcomes—either of a cosmology that crunches (usual wormholes), or of a cosmology that is expanding (wineglass and quasi-oscillatory wormholes). In particular, wineglass wormholes Betzios:2024oli ; Betzios:2024zhf can become the leading connected contribution for sufficiently small values of the electromagnetic field at the UV boundary, and under analytic continuation they naturally give rise to expanding cosmological spacetimes with an early inflationary phase.
@@
The wineglass AdS wormholes were first devised Betzios:2024oli ; Betzios:2024zhf as a mean to surpass the probabilistic issues inflicting the Hartle-Hawking no-boundary proposal (see Lehners:2023yrj ; Maldacena:2024uhs for reviews) and the resulting trivial (one-dimensional) Hilbert space for the closed cosmologies. For these reasons, a clear holographic understanding of closed Universes within the no-boundary proposal has been elusive so far (see though Liu:2025cml ; Antonini:2025ioh ; Harlow:2026hky ; Abdalla:2026mxn ; Nomura:2026igt ; Zhao:2026mpl for some recent works with a perspective emphasizing the role of observers in this setting). On the other hand, the AdS wormhole proposal of Betzios:2024oli ; Betzios:2024zhf can give rise to a non-trivial Hilbert space for (expanding/inflating) cosmologies, through the field theory dual living on the AdS asymptotic boundary/ies, and realised semi-classically via the different competing saddles and asymptotic boundary conditions for the various bulk fields.
@@
We conclude this manuscript by presenting a list of topics that deserve further analysis. First, it would be interesting to study ours or similar models for more general parameters (even numerically), to explore and carve out the parameter regimes in which wineglass or other oscillatory wormholes that can give rise to an early inflationary cosmological outcome are the dominant saddles of the gravitational path integral. As a variant of this idea, magnetic types of wormholes can also be found in braneworld scenarios Antonini:2024bbm . It would be interesting to consider more general types of wormholes in such scenarios and verify whether they can lead to inflating braneworld cosmologies.
@@
The family of models that we study in this work constitutes the first instance where a class of models that admits such an intricate variety of wormholes and disconnected saddles, is analysed and discussed in the literature in higher dimensions. In two spacetime dimensions there do exist simple models that contain a similar plethora of different kinds of gravitational saddles (i.e. the “gravitational Sine-Gordon model” Betzios:2025eev ) crucially relying again on a scalar matter potential that can take both positive and negative values. Since these two dimensional models also admit a matrix model description, it is very important to analyse them in great detail, to learn about the non-perturbative and microscopic physics in such rich instances of the gravitational path integral.
@@
At the same time, it is also of utmost importance to perform a careful study of cosmological correlators using these saddles, to understand the similarities and differences with the no-boundary proposal and ascertain their potential role in a more phenomenological setting Betzios2025toappear .
@@
Finally, since the class of models that we consider in this work combines elements that are generic in flux compactifications in string theory (a non-trivial scalar potential containing both AdS/negative and dS/positive local minima as well as local positive maxima) together with the presence of non-trivial fluxes, we believe that it is worthwhile to extend our study to models that arise in more realistic string compactifications. In particular the recent work AbdusSalam:2025twp discusses such cases of coexisting vacua, which can also realise our proposal within the context of supergravity.

### v2

@@ 4.4.2 The oscillatory wormhole E/M gauge field background
The “Centaur” geometries of Anninos and Hofman Anninos and Hofman (2018), can in a sense be thought of as a finely tuned case of the wineglass wormholes, when the running scalar reaches a local minimum/maximum of the scalar potential with positive value and stops completely there. On the one hand, it seems that it is always possible to achieve this, for example if the potential $V(\phi)$ becomes exactly flat after a certain value of $\phi$. Unfortunately the presence of the magnetic flux/radiation, which is very important for supporting the throat at its minimum value $a_{\rm min}(\tau_{\rm min})$, prohibits also the smooth closure of the geometry at other Euclidean times.
@@ 5.1 Dirichlet for the scalar field/Dirichlet for the gauge field
Thus, we expect $\Delta\mathcal{S}_{D-C}\underset{H_{\rm UV}\to\infty}{\sim}S_{E,\rm rad}^{D}>0$ for large $H_{\rm UV}$, i.e., the simple wormhole being more dominant in the gravitational path integral. The complete numerical results are shown fig. 8  [^24] In the right panel, we also show the results imposing Neumann BCs for the gauge field. The results are qualitatively similar.. In a certain regime there exist two wormholes: one is always subdominant relative to the disconnected geometry and ceases to exist when reaching the critical value $(\kappa|\tilde{V}_{\rm AdS}|)^{1/2}H_{\rm UV}=2\sqrt{2}$. This is the bound found earlier for the coexistence of two simple wormholes, see fig. 2. Keeping the dominant saddle between these two wormholes, we observe an exchange of dominance between the disconnected geometry and the wormhole [^25] This was also observed in two dimensional string theory models that exhibit BH/wormhole transitions (two disconnected cigars vs. a simple Euclidean wormhole) Betzios et al. (2024). revealing a first order phase transition similar to the Hawking-Page transition at $(\kappa|\tilde{V}_{\rm AdS}|)^{1/2}H_{\rm UV}\simeq 2.708$ Hawking and Page (1983).
@@ 5.2 Neumann for the scalar field/Dirichlet for the gauge field
The second line corresponds precisely to the counterterm and the boundary term. The counterterm is computed in Appendix A.1, and the boundary term vanishes on-shell. This is because, when $c_{\rm AdS}\neq-2$, the mode $\tilde{\phi}_{+}$ (proportional to the source using the Neumann boundary condition) is zero. Note that, in contrast to the preceding section, $c_{\rm AdS}$ is now related to a VEV and thus free to vary, resulting in a four-parameter family which we parametrize in terms of $H_{\rm UV},\,\tilde{V}_{\rm AdS},\,A_{\rm dS},\,A_{\rm AdS}$. After sending the cut-offs to infinity, the difference between the on-shell actions remains finite
@@ 6 Discussion
    Semiclassical saddle competition in the gravitational path integral and Cosmological ramifications: By explicitly evaluating the on-shell action differences, we identified the parameter regions in which different saddles dominate the gravitational path integral. As we explain in section 2, the connected two boundary saddles can be used to estimate ratios of probabilities for different cosmological outcomes—either of a cosmology that crunches (usual wormholes), or of a cosmology that is expanding (wineglass and quasi-oscillatory wormholes). In particular, wineglass wormholes Betzios and Papadoulaki (2024a); Betzios et al. (2025) can become the leading connected contribution for sufficiently small values of the electromagnetic field at the UV boundary, and under analytic continuation they naturally give rise to expanding cosmological spacetimes with an early inflationary phase.
@@
From the class of backgrounds that we constructed, two deserve some further discussion. The wineglass AdS wormholes were first devised Betzios and Papadoulaki (2024a); Betzios et al. (2025) as a mean to surpass the probabilistic issues inflicting the Hartle-Hawking no-boundary proposal (see Lehners (2023); Maldacena (2024) for reviews) and the resulting trivial (one-dimensional) Hilbert space for the closed cosmologies. For these reasons, a clear holographic understanding of closed Universes within the no-boundary proposal has been elusive so far (see though Liu (2025); Antonini et al. (2025); Harlow (2026); Abdalla et al. (2026); Nomura and Ugajin (2026); Zhao (2026) for some recent works with a perspective emphasizing the role of observers in this setting). On the other hand, the AdS wormhole proposal of Betzios and Papadoulaki (2024a); Betzios et al. (2025) can give rise to a non-trivial Hilbert space for (expanding/inflating) cosmologies, through the field theory dual living on the AdS asymptotic boundary/ies, and realised semi-classically via the different competing saddles and asymptotic boundary conditions for the various bulk fields.
Oscillatory wormholes were first constructed in Halliwell and Myers (1989) and recently discussed in Aguilar-Gutierrez et al. (2023) using models containing axionic fields. The axion wormhole models seem to suffer from the issue that the flat direction of the (classically) non-existent axion potential (one just considers a fixed positive cosmological constant in these models), leads to the potential presence of an unbound infinite number of oscillations for such solutions and to a badly divergent path integral (an avatar of the notorious conformal mode problem of the Euclidean gravitational path integral). As we describe in section 4.4, a breaking of the shift symmetry of the potential precludes this issue, giving rise to an upper bound on the number of oscillations. While in the present work we studied this mechanism in the context of Einstein-Scalar-Radiation models, it is straightforward to apply it also to the context of axionic wormholes. In particular it is well known that non-perturbative instanton or gravitational effects (see Hebecker et al. (2018) for a review) lead to an effective potential for the axionic field $\alpha$ of the form
$V_{\rm eff}(\alpha)=\sum_{n}e^{-S_{n}}\cos\left(\frac{n\alpha}{f}+\delta_{n}\right)\,$, (118)
where $S_{n}$ is the on-shell action of the instanton-like configuration, $f$ is the axion coupling strength ($\mathcal{L}(\alpha)\sim f^{2}(\partial\alpha)^{2}$) and $\delta_{n}$ are model dependent coefficients (incommensurate). The presence of scalar fields will render the effective potential an even more complicated function. Thus the issue of the infinite number of oscillations is completely resolved in any realistic physical model, by lifting the naive flat direction(s) into a non-trivial potential.
@@
We conclude this manuscript by presenting a list of topics that deserve further analysis. First, it would be interesting to study ours or similar models for more general parameters (even numerically), to explore and carve out the parameter regimes in which wineglass or other oscillatory wormholes that can give rise to an early inflationary cosmological outcome are the dominant saddles of the gravitational path integral. As a variant of this idea, magnetic types of wormholes can also be found in braneworld scenarios Antonini and Bariuan (2024). It would be interesting to consider more general types of wormholes in such scenarios and verify whether they can lead to inflating braneworld cosmologies.
@@
The family of models that we study in this work constitutes the first instance where a class of models that admits such an intricate variety of wormholes and disconnected saddles, is analysed and discussed in the literature in higher dimensions. In two spacetime dimensions there do exist simple models that contain a similar plethora of different kinds of gravitational saddles (i.e. the “gravitational Sine-Gordon model” Betzios (2026)) crucially relying again on a scalar matter potential that can take both positive and negative values. Since these two dimensional models also admit a matrix model description, it is very important to analyse them in great detail, to learn about the non-perturbative and microscopic physics in such rich instances of the gravitational path integral.
@@
At the same time, it is also of utmost importance to perform a careful study of cosmological correlators using these saddles, to understand the similarities and differences with the no-boundary proposal and ascertain their potential role in a more phenomenological setting Betzios et al. (To appear).
@@
Finally, since the class of models that we consider in this work combines elements that are generic in flux compactifications in string theory (a non-trivial scalar potential containing both AdS/negative and dS/positive local minima as well as local positive maxima) together with the presence of non-trivial fluxes, we believe that it is worthwhile to extend our study to models that arise in more realistic string compactifications. In particular the recent work AbdusSalam et al. (2026) discusses such cases of coexisting vacua, which can also realise our proposal within the context of supergravity.